# Probabilistically-sound beam search with masked language models

### v1

@@ Abstract
Beam search with masked language models (MLMs) is challenging in part because joint probability distributions over sequences are not readily available, unlike for autoregressive models. Nevertheless, estimating such distributions has applications in many domains, including protein engineering and ancient text restoration. We present probabilistically-sound methods for beam search with MLMs. First, we clarify the conditions under which it is theoretically sound to perform text infilling with MLMs using standard beam search. When these conditions fail, we provide a probabilistically-sound modification with no additional computational complexity and demonstrate that it is superior to the aforementioned beam search in the expected conditions. We then present empirical results comparing several infilling approaches with MLMs across several domains.
@@ 1 Introduction
Autoregressive language models (LMs) have showcased exceptional ability across many tasks. Yet, in specific contexts where bidirectionality is crucial, such as protein language modeling and ancient text restoration, masked language models (MLMs) remain prevalent. However, MLMs still face a significant challenge in these settings: while MLMs learn conditional distributions of single tokens, applications to the aforementioned domains often require knowledge of the probability of multiple tokens jointly.
The key challenge lies in computing the joint distribution $p(\mathbf{x})$ over sequences $\mathbf{x}$ given only the MLM-learned conditionals $p(x_{i}|\mathbf{x}_{-i})$, where $\mathbf{x}_{-i}$ denotes the context sequence $\mathbf{x}$ with the entry at index $i$ removed. The Hammersley-Clifford-Besag (HCB) theorem yields a direct algebraic construction of a joint distribution $p(\mathbf{x})$ which is valid in the case when the MLM-learned distributions are compatible—that is, there exists a joint distribution which factors into the conditionals exactly Hennigen and Kim (2023).
@@
In practical applications, a heuristic approximation based on the chain rule of autoregressive LMs is used for infilling with MLMs (Ruffolo and Madani, 2024; Shen et al., 2020; Assael et al., 2022; Cowen-Breen et al., 2023b):
@@
Our first contribution is demonstrating that Equation 1 is valid if and only if a conditional independence assumption about the MLM-learned conditionals is satisfied and providing the conditions under which this assumption holds, assuming compatibility. We state these conditions in Theorem 1:
@@
We hypothesize that there may be two regimes as a consequence of Theorem 1: a regime where training loss is small enough that the heuristic approach of Equation 1 may be reasonable, and a regime where the conditional independence assumption may not hold, in which another approach is needed.
Our second contribution is providing a modification of Equation 1 which holds for the second, possibly under-trained regime. Based on the HCB theorem, this modification relaxes the conditional independence assumption by including an adjustment term to correct for possible dependencies, and it requires no additional forward passes through the model.
Evidently, not all MLMs of interest lie in the regime where Equation 1 is reasonable: we find empirically that our modification outperforms standard beam search for certain models, including BERT-base. In developing RoBERTa, Liu et al. (2019) suggest that BERT was under-trained, which aligns with our experimental observation that the conditional independence assumption seems more reasonable for RoBERTa than for BERT.
@@ 2 Text infilling
Empirically, we evaluate Equation 1 and our modification of it on the task of text infilling – predicting a missing span of text given its context – which is well-motivated in the domains of protein language modeling and ancient text restoration Zhu et al. (2019).
Much existing work in text infilling focuses on developing custom architectures and training from scratch Sun et al. (2017); Ippolito et al. (2019); Shen et al. (2020); Donahue et al. (2020). In some domains, however, pretrained MLMs may be available in instances where limits on available data or compute prohibit training new text infilling models. Manuscript restoration is a prototypical example: training data is often restricted by copyright (Graziosi et al., 2023), and full trainings can be computationally expensive, yet various pretrained MLMs for ancient languages are available Bamman and Burns (2020); Assael et al. (2022); Cowen-Breen et al. (2023a); Riemenschneider and Frank (2023). Another example is protein language models (PLMs): the computational footprint of fine-tuning PLMs becomes a barrier for many research groups Hu et al. (2021); Sledzieski et al. (2023). Despite the challenges posed by data and compute, effective infilling remains an important task for ancient text restoration and protein engineering.
Here, we examine the capabilities of MLMs to infill directly, primarily by infilling tokens sequentially through an adaptation of beam search to MLMs, although we compare additionally to other sampling schemes, such as nucleus sampling and sampling with temperature.
It should be noted that our methods only apply to settings where the number of missing tokens is fixed, as opposed to the more general setting addressed by Shen et al. (2020), among others. That said, fixing the number of missing tokens is reasonable in the case of infilling damaged texts or missing amino acids. [^1] In the case of damaged inscriptions, domain experts posit an estimated number of missing characters based on physical distance Bruun and Edmondson (2014).
@@ 3.1 Beam search
Beam search is a form of decoding which incrementally adds successive tokens to a set of candidate token sequences, maintaining a fixed number of candidate sequences $\mathbf{x}=(x_{1},\ldots,x_{n})$ of highest joint probability $p(\mathbf{x})$. For autoregressive models, the chain rule [^2] In what follows, we generally omit right-context tokens from $p(\cdot|\cdot)$ for readability, but we note that modified forms of all of the following equations hold when $p$ is further conditioned on right-context.,
@@
allows for the implementation of beam search shown in Algorithm 1.
Initialize $S=\{(0,\emptyset)\}$
for $i\in\{1,\ldots,n\}$ do
     for $(\ell,(x_{1},\ldots,x_{i-1}))\in S$ do
         $f(\cdot)\leftarrow\log p(\cdot|{x_{1},\ldots,x_{i-1}})$ (1 forward pass.33footnotemark: 3)
         Append to $S$: $(\ell+f(x),(x_{1},\ldots,x))$ for every $x$.
     end for
[…]
end for
return S
@@ 3.3 Constructing joint distributions from conditionals
$\begin{split}\log p(\mathbf{x})\sim\sum_{i=1}^{n}&\log p(x_{i}|\mathbf{x}_{:i},\mathbf{y}_{i:})\\ -&\log p(y_{i}|\mathbf{x}_{:i},\mathbf{y}_{i:})\end{split}$ (5)
where $\sim$ indicates equality up to addition of a constant in $\mathbf{x}$. Following Hennigen and Kim (2023), we refer to $\mathbf{y}$ as the pivot. When the conditional distributions are compatible, Equation 5 should yield orderings of sequences $\mathbf{x}$ by their probabilities $p(\mathbf{x})$ in a manner which is consistent across choice of pivots.
@@ 4 Methods
Our primary observation is that the HCB theorem (Theorem 2) yields a straightforward correction to the standard beam search induced by Equation 4 which incurs almost no additional computational cost. [^7] Code for our experiments can be found at https://github.com/rcalef/hcb_infilling Although the conditional distributions learned by BERT are not exactly compatible, they are empirically compatible enough to improve the accuracy of beam search in certain instances. Our hope is that these methods will be useful for infilling when the implementation of the AG construction is intractable.
@@
$\begin{split}\log p(\mathbf{x})\sim\sum_{i=1}^{n}&\log p(x_{i}|\mathbf{x}_{:i},\mathbf{[M]}_{i:})\\ -&\log p([\operatorname{M}]|\mathbf{x}_{:i},\mathbf{[M]}_{i:})\end{split}$ (6)
where $\sim$ again indicates equality up to addition of a constant. This correction term guarantees probabilistic soundness of the type ensured by Equation 2 and requires no additional forward passes to compute, as the tensor $p(\cdot|\mathbf{x}_{:i},\mathbf{[M]}_{i:})$ is computed with a single forward pass. Although the value of $p([\operatorname{M}]|\mathbf{x}_{:i},\mathbf{[M]}_{i:})$ does not have an immediately obvious intuitive meaning, this quantity empirically contains enough signal [^8] In particular, we find that it depends strongly on the context, $\mathbf{x}_{:i}$. to demonstrate improvements over baselines and several ablations, as we demonstrate in Section 6.
@@ 4.1 Choosing a pivot
The MLM training procedure makes one particular choice of pivot especially in-distribution for any context: the sequence of mask tokens $\mathbf{y}=([\operatorname{M}],\ldots,[\operatorname{M}])$. Throughout MLM training, mask locations are re-randomized, meaning that the MLM is likely to encounter sequences of masks as context in many positions across all examples in the train set. The probability $p([\operatorname{M}]|\mathbf{x}_{<i},\mathbf{y}_{>i})$ decreases drastically during training, as $[\operatorname{M}]$ never occurs as a label. [^9] See Figure 7 for how $p([\operatorname{M}|\cdot])$ changes during BERT training. Note that while $[\operatorname{M}]$ is never seen as a label during training, it is still awarded positive probability on account of the softmax in the final layer of BERT. Nonetheless, using this quantity improves infilling accuracy across different models in various domains. We perform ablation studies to verify this probability captures genuine information about the context, rather than random noise.
5 Experimental setup
5.1 Models
5.2 Metrics
5.3 Datasets
In English, we perform infilling experiments on three datasets: the Brown corpus Nelson Francis and Kucera (1979), the Stanford Natural Language Inference dataset (SNLI) Bowman et al. (2015), and the extreme summarization dataset (XSUM) Narayan et al. (2018). This allows us to additionally test varying amounts of context: the datasets have an average of 452, 19, and 28 tokens of context per example, respectively. For ancient language models, we perform infilling experiments on the Packhard Humanities Institute’s database of ancient Greek inscriptions, [^11] https://inscriptions.packhum.org/ collected by Sommerschield* et al. (2021). For protein language models, we perform infilling experiments on a subset of protein sequences from UniProt UniProt (2008).
5.4 Baselines
5.5 Ablations
To test that the estimated probability of the mask token $p([\operatorname{M}]|\cdot)$ is not simply random noise and leads to improved beam search in several settings, we perform two ablations.
Ablation 1 (Context Scramble):
Ablation 2 (Random Token Swap):
6 Results
6.1 English language models
In our English language experiments with BERT-base, we observe a consistent relative ranking of methods: HCB Best-to-Worst $>$ HCB Left-to-Right $>$ Standard Best-to-Worst $>$ Standard Left-to-Right. This ranking persists across metrics (top-$k$ accuracy and BLEU; Table 3), number of missing tokens (2 through 5; Table 2), datasets (Brown, SNLI, XSUM; Table 1), and beam sizes (5 and 20; Figure 2).
Model BERT-base RoBERTa
$\mathbb{B}$ HCB Standard 11.43 25.29
10.63 25.81
$\mathbb{S}$ HCB Standard 7.49 13.66
7.08 14.56
$\mathbb{X}$ HCB Standard 9.86 23.21
9.18 24.52
Table 1: Comparison of HCB vs. standard beam search on the Brown $\mathbb{B}$, SNLI $\mathbb{S}$ and XSUM $\mathbb{X}$ datasets, across models. We only show best-to-worst, since left-to-right was strictly worse in each case.
We observe that BERTScore values are largely consistent across beam search methods (Table 3), likely due to the relatively small number of tokens being infilled in each example, but HCB consistently outperforms standard beam search.
Note that standard left-to-right beam search outperforms our nucleus sampling-beam search hybrid, consistent with the findings of Shaham and Levy (2022), as well as all other sampling-based methods.
Across models, we note that the various beam search methods rank similarly when using BERT-large and DistilBERT, but RoBERTa stands out as a case where standard beam search outperforms HCB beam search. For a given model, we also find that the ranking of methods is consistent across values of $k$ for top-$k$ accuracy (Figure 2).
@@ 6.1 English language models
Table 2: Top-$k$ accuracy at infilling consecutive missing tokens on 100K examples from Brown corpus, using BERT-base with beam size 5 and number of missing tokens ranging from 2 to 5.
@@
Table 3: BLEU score and BERTScore at infilling consecutive missing tokens on 100K examples from Brown corpus, using BERT-base, with beam size 5.
Figure 2: Performance of HCB beam search versus standard beam search methods across three English MLMs and DistilBERT, with beam size 20. Note that RoBERTa appears to be suffering from the beam search curse, as noted by Meister et al. (2020).
@@ 6.2 Domain-specific language models
Protein sequences. In experiments infilling between two and five missing amino acids in a protein sequence using ESM2, we find that HCB beam search shows comparable, albeit lower performance to standard beam search (Appendix B, Table 9). Thus, it appears that for protein sequences, the conditional independence assumption may hold better than for human language, leading to relatively strong standard beam search performance; this is further explored in Section 7.
@@ 6.3 Pivot design
To explore how performance of HCB beam search depends on the choice of pivot, we extensively test all pivots for seven-character infills with a beam size of 20 for Ithaca. Testing every possible pivot choice is made possible by Ithaca’s small 34-token vocabulary. With over 800 trials per pivot, we observe that the special token “-”, which corresponds to a missing character, is clearly the best performing pivot choice, achieving accuracy similar to standard beam search. Two other special tokens “#” and “$<$” perform about $20\%$ worse as HCB pivots, and the 31 remaining tokens perform over $50\%$ worse than standard beam search (Figure 3). We also exhaustively test all pivots for ESM2, observing that the best choice of pivot appears to be the special token “[CLS]” (Appendix B, Figure 6), though this optimal pivot choice is much less clear than for Ithaca.
Notably, Ithaca’s architecture does not include the mask token in its output probability distribution; thus, the missing character token, which is scattered randomly throughout train and test examples, is the token which remains most in-distribution when injected as a pivot into random positions. Similarly, for ESM2, we observe that some special tokens perform noticeably better as pivots than regular tokens. These results could suggest that special tokens, which do not add disruptive context, are better candidates for pivots.

### v3

# Towards Probabilistically-Sound Beam Search with Masked Language Models

@@ Abstract
Beam search with masked language models (MLMs) is challenging in part because joint probability distributions over sequences are not readily available, unlike for autoregressive models. However, estimating such distributions has important domain-specific applications such as ancient text restoration and protein engineering. Here we present probabilistically-sound methods for beam search with MLMs. First, we clarify the conditions under which it is theoretically sound to perform text infilling with MLMs using standard beam search. When these conditions fail, we provide a probabilistically-sound inference time modification with no additional computational complexity and demonstrate that it is superior to the aforementioned beam search in the expected conditions. We then present empirical results comparing several infilling approaches with MLMs across several domains. Notably, our method probes the inductive biases of MLMs and explores the surprising contextual sensitivity of mask tokens for text infilling.
@@ 1 Introduction
Autoregressive language models (LMs) have demonstrated success in many tasks. Yet in specific contexts where bidirectionality is crucial, such as ancient text restoration and protein engineering, masked language models (MLMs) are most prevalent. Notably, MLMs still face a significant challenge in these settings: while MLMs learn conditional distributions of single tokens, applications to the aforementioned domains often require knowledge of the probability of multiple tokens jointly.
The key challenge lies in computing the joint distribution $p(\mathbf{x})$ over sequences $\mathbf{x}$ given only the MLM-learned conditionals $p_{i}(x_{i}|\mathbf{x}_{-i})$, where $\mathbf{x}_{-i}$ denotes the context sequence $\mathbf{x}$ with the entry at index $i$ removed. The Hammersley-Clifford-Besag (HCB) theorem (Besag, 1974) yields a direct algebraic construction of a joint distribution $p(\mathbf{x})$ which is valid in the case when the MLM-learned distributions are compatible—that is, there exists a joint distribution which factors into the conditionals exactly Hennigen and Kim (2023).
@@
In practical applications, a heuristic approximation based on the chain rule of autoregressive LMs is used for infilling with MLMs (Ruffolo and Madani, 2024; Shen et al., 2020; Assael et al., 2022; Cowen-Breen et al., 2023b; Wang and Cho, 2019):
@@
In this work, we demonstrate that Equation 1 is valid if and only if a conditional independence assumption about the MLM-learned conditionals is satisfied and provide the conditions under which this assumption holds, assuming compatibility. We state these conditions in Theorem 1:
@@
We hypothesize that there may be two regimes as a consequence of Theorem 1: a regime where training loss is small enough that the heuristic approach of Equation 1 may be reasonable, and a regime where the conditional independence assumption may not hold, in which another approach is needed. In this work, we provide a modification of Equation 1 based on the HCB theorem in the second regime, which relaxes the conditional independence assumption by including an adjustment term to correct for possible dependencies and requires no additional forward passes through the model. We present this modification, which we call HCB beam search, and the resulting theoretical and empirical contributions to zero-shot text infilling.
Theoretical contributions.
Previous work has evaluated text scoring rather than text infilling, considering proxies for joint distributions to evaluate the probabilities of pairs of tokens and computing the perplexities of these joint distributions Hennigen and Kim (2023). A key insight of our work is that the HCB construction is the only such proxy which decomposes in a way suitable for beam search. It is also particularly useful in low-resource settings, as it can be implemented with no additional computational requirements, allowing for the adaptation of generalist MLMs for the text infilling task. We present a theorem classifying the conditions under which the compatibility of learned distributions hold and showing that standard beam search with MLMs is justified precisely when the conditions of this theorem hold.
Empirical contributions.
[…]
@@ 2 Text Infilling
We experimentally evaluate Equation 1 and our modification of it via text infilling, the task of predicting a missing span of text given its context. This task is well-motivated in the domains of protein language modeling and ancient text restoration Zhu et al. (2019).
Much existing work on text infilling focuses on developing custom architectures and training from scratch Sun et al. (2017); Ippolito et al. (2019); Shen et al. (2020); Donahue et al. (2020); He et al. (2023). In some low-resource domains, however, pretrained MLMs may be available when limitations on available data or compute prohibit training new text infilling models. Manuscript restoration is a prototypical example: training data is often restricted by copyright (Graziosi et al., 2023), and full trainings can be computationally expensive, yet various pretrained MLMs for ancient languages are available Bamman and Burns (2020); Cowen-Breen et al. (2023a); Riemenschneider and Frank (2023). Another example is protein language models (PLMs): the computational footprint of fine-tuning PLMs becomes a barrier for many research groups Hu et al. (2021); Sledzieski et al. (2023). Despite the challenges posed by data and compute, effective infilling remains an important task for ancient text restoration and protein engineering.
@@ 3.1 Beam search
Beam search is a form of decoding which incrementally adds successive tokens to a set of candidate token sequences, maintaining a fixed number of candidate sequences $\mathbf{x}=(x_{1},\ldots,x_{n})$ of highest joint probability $p(\mathbf{x})$. For autoregressive models, the chain rule, [^1] In what follows, we generally omit right-context tokens from $p(\cdot|\cdot)$ for readability, but we note that modified forms of all of the following equations hold when $p$ is further conditioned on right-context.
@@
allows for the well-known autoregressive beam search algorithm (Appendix B, Algorithm 2).
@@ 3.2 Challenges with MLM beam search
Scoring functions $f$ for various beam search implementations: [^4] The notation $\mathbf{[M]}_{i:k}$ indicates that [mask] tokens occupy the indices from $i$ to $k$. Unless otherwise specified, the left-most mask index ($i$) is the one at which the probability $p(\cdot|\mathbf{[M]}_{i:k})$ will be evaluated.
Standard : $\text{Standard :\quad\quad}f(\cdot)$ $=\log p(\cdot|\mathbf{x}_{:i},\mathbf{[M]}_{i:k},\mathbf{x}_{k:})$
[…]
@@ 3.3 Constructing joint distributions from conditionals
$\begin{split}\log p(\mathbf{x})\sim\sum_{i=1}^{n}&\log p(x_{i}|\mathbf{x}_{:i},\mathbf{y}_{i+1:})\\ -&\log p(y_{i}|\mathbf{x}_{:i},\mathbf{y}_{i+1:})\end{split}$ (5)
where $\sim$ indicates equality up to addition of a constant $\mathbf{x}$ (equal to $\log p(\mathbf{y})$). Following Hennigen and Kim (2023), we refer to $\mathbf{y}$ as the pivot. When the conditional distributions are compatible, Equation 5 should yield orderings of sequences $\mathbf{x}$ by their probabilities $p(\mathbf{x})$ in a manner which is consistent across choice of pivots.
@@ 4 Methods
Our primary observation is that the HCB theorem (Theorem 2) yields a straightforward correction to the standard beam search induced by Equation 4 which incurs almost no additional computational cost. [^5] Code for our experiments can be found at https://github.com/rcalef/hcb_infilling Although the conditional distributions learned by BERT are not exactly compatible, they are empirically compatible enough to improve the accuracy of beam search in certain instances. [^6] It is worth noting that standard beam search assumes compatibility to the same extent that HCB does. Thus, these methods can be useful for infilling when the implementation of the AG construction is intractable.
@@
$\begin{split}\log p(\mathbf{x}_{j:k}|\mathbf{x}_{:j},\mathbf{x}_{k:})\sim\sum_{i=j}^{k}&\log p(x_{i}|\mathbf{x}_{:i},\mathbf{[M]}_{i:k},\mathbf{x}_{k:})\\ -&\log p([\operatorname{M}]|\mathbf{x}_{:i},\mathbf{[M]}_{i:k},\mathbf{x}_{k:})\end{split}$ (6)
where $\sim$ again indicates equality up to addition of a constant in the tokens-to-be-infilled $\mathbf{x}_{j:k}$, equal to $\log p([\mathbf{M}]_{i:k}|\mathbf{x}_{:j},\mathbf{x}_{k:})$. This correction term guarantees probabilistic soundness of the type ensured by Equation 2 and requires no additional forward passes to compute, as the tensor $p(\cdot|\mathbf{x}_{:i},\mathbf{[M]}_{i:})$ is computed with a single forward pass. Although the value of $p([\operatorname{M}]|\mathbf{x}_{:i},\mathbf{[M]}_{i:})$ does not have an immediately obvious intuitive meaning, Ablation 1 and Ablation 2 show that it depends strongly on the context $\mathbf{x}_{:i}$ and can be used to improve over baselines.
@@ 4.1 Choosing a pivot
The MLM training procedure makes one particular choice of pivot especially in-distribution for any context: the sequence of mask tokens $\mathbf{y}=([\operatorname{M}],\ldots,[\operatorname{M}])$. Throughout MLM training, mask locations are re-randomized, meaning that the MLM is likely to encounter sequences of masks as context in many positions across all examples in the train set. The probability $p([\operatorname{M}]|\mathbf{x}_{<i},\mathbf{y}_{>i})$ decreases drastically during training, as $[\operatorname{M}]$ never occurs as a label. [^7] See Figure 6 for how $p([\operatorname{M}|\cdot])$ changes during BERT training. Note that while $[\operatorname{M}]$ is never seen as a label during training, it is still awarded positive probability on account of the softmax in the final layer of BERT. Nonetheless, using this quantity improves infilling accuracy across different models in various domains. We perform ablation studies to verify that this probability captures genuine information about the context, rather than random noise.
Number of missing tokens 2 3 4 5 2 3 4 5
@@
Table 1: Top-$k$ accuracy, BLEU score, and BERTScore at infilling consecutive missing tokens on 100K examples from Brown corpus, using BERT-base with beam size 5 and number of missing tokens ranging from 2 to 5.
5 Experimental Setup
5.1 Models
5.2 Metrics
5.3 Datasets
In English, we perform infilling experiments on three datasets: the Brown corpus Nelson Francis and Kucera (1979), the Stanford Natural Language Inference dataset (SNLI) Bowman et al. (2015), and the extreme summarization dataset (XSUM) Narayan et al. (2018). This allows us to additionally test varying amounts of context: the datasets have an average of 452, 19, and 28 tokens of context per example, respectively. For ancient language models, we perform infilling experiments on the Packhard Humanities Institute’s database of ancient Greek inscriptions, collected by Sommerschield* et al. (2021). For protein language models, we perform infilling experiments on a subset of protein sequences from UniProt UniProt (2008).
5.4 Baselines
5.5 Ablations
To test that the estimated probability of the mask token $p([\operatorname{M}]|\cdot)$ is not simply random noise, as we found a pivot of mask tokens is crucial for the success of HCB beam search, we perform two ablations.
Ablation 1 (Context Scramble):
Ablation 2 (Random Token Swap):
6 Results
6.1 English language models
In our English language experiments with BERT-base, we observe a consistent relative ranking of methods: HCB Best-to-Worst $>$ HCB Left-to-Right $>$ Standard Best-to-Worst $>$ Standard Left-to-Right. Table 1 shows that this ranking persists across metrics (top-$k$ accuracy, BLEU, BERTScore F1) and number of missing tokens (2 through 5). In Table 2 and Figure 4, we see the ranking additionally persists across datasets (Brown, SNLI, XSUM) and beam sizes (5 and 20), respectively.
We observe that BERTScore values are largely consistent across beam search methods (Table 3), likely due to the relatively small number of tokens being infilled in each example, but HCB consistently outperforms standard beam search. We note too that standard left-to-right beam search outperforms our nucleus sampling-beam search hybrid, consistent with the findings of Shaham and Levy (2022), as well as all other sampling-based methods.
Across models we note that the various beam search methods rank similarly when using BERT-large and DistilBERT, though RoBERTa stands out as a case where standard beam search outperforms HCB beam search. For a given model, we also find that the ranking of methods is consistent across values of $k$ for top-$k$ accuracy (Appendix C, Figure 4).
Model BERT-base RoBERTa
$\mathbb{B}$ HCB Standard 11.43 25.29
10.63 25.81
$\mathbb{S}$ HCB Standard 7.49 13.66
7.08 14.56
$\mathbb{X}$ HCB Standard 9.86 23.21
9.18 24.52
Table 2: Comparison of top-1 accuracy for HCB vs. standard beam search on the Brown $\mathbb{B}$, SNLI $\mathbb{S}$ and XSUM $\mathbb{X}$ datasets, across models. We only show best-to-worst, since left-to-right was strictly worse in each case.
@@ 6.2 Domain-specific language models
Protein sequences. In experiments infilling between two and five missing amino acids in a protein sequence using ESM2, we find that HCB beam search shows comparable, albeit lower performance to standard beam search (Appendix C, Table 6). Thus, it appears that for protein sequences, the conditional independence assumption may hold better than for human language, leading to relatively strong standard beam search performance; this is further explored in Section 7.
@@ 6.3 Pivot design
To explore how performance of HCB beam search depends on the choice of pivot, we extensively test all pivots for seven-character infills with a beam size of 20 for Ithaca. Testing every possible pivot choice is made possible by Ithaca’s small 34-token vocabulary. With over 800 trials per pivot, we observe that the special token “-”, which corresponds to a missing character, is clearly the best performing pivot choice, achieving accuracy similar to standard beam search. Two other special tokens “#” and “$<$” perform about $20\%$ worse as HCB pivots, and the 31 remaining tokens perform over $50\%$ worse than standard beam search (Figure 2). We also exhaustively test all pivots for ESM2, observing that the best choice of pivot appears to be the special token “[CLS]” (Appendix C, Figure 5), though this optimal pivot choice is much less clear than for Ithaca.
Notably, Ithaca’s architecture does not include the mask token in its output probability distribution; thus, the missing character token, which is scattered randomly throughout train and test examples, is the token which remains most in-distribution when injected as a pivot into random positions. Similarly, for ESM2, we observe that some special tokens perform noticeably better as pivots than regular tokens. These results suggest that special tokens, which do not add disruptive context, could be better candidates for pivots.